OpenStudy (princeevee):

solve x-9=-15 -24 24 6 -6

OpenStudy (just_one_last_goodbye):

Need help?

OpenStudy (bossvideogamer21):

-6

OpenStudy (bossvideogamer21):

is the answer

OpenStudy (just_one_last_goodbye):

Problem: x-9=-15 Steps (Add 9 to both sides): x - 9 + 9 = -15 + 9 Solve for x: answer is ?

OpenStudy (just_one_last_goodbye):

Please No direct answer @Bossvideogamer21 read the CoC http://openstudy.com/code-of-conduct

OpenStudy (bossvideogamer21):

oh i am so sorry i didnt know

OpenStudy (princeevee):

thanks

OpenStudy (just_one_last_goodbye):

:/ ya alright...

OpenStudy (bossvideogamer21):

i read it @just_one_last_goodbye

OpenStudy (just_one_last_goodbye):

So please follow it for now on ^_^

OpenStudy (bossvideogamer21):

ok